# ForgetMark: Stealthy Fingerprint Embedding via Targeted Unlearning in Language Models

### v1

@@ Abstract
footnotetext: *Equal contribution. Corresponding author.
@@ 1 Introduction
Large language models (LLMs) and related paradigms such as vision–language models [50, 48], autonomous agents [47, 49], and AI systems for security applications [17, 18, 37, 32, 43] have become central to modern machine learning research and deployment, yet their creation depends on substantial proprietary data, compute, and engineering. Illicit model exfiltration and unauthorized fine-tuning have lowered the barrier to redistributing derivatives, creating urgent needs for reliable provenance auditing and copyright protection.
Model fingerprinting [36] has emerged as a practical remedy. Existing approaches fall into two families: intrinsic fingerprints, which rely on internal parameters or representations [5, 44, 42, 45], and invasive (backdoor-based) fingerprints, which implant a trigger–response mechanism detectable under black-box access [28, 41, 34, 4, 38]. Intrinsic methods often require access to weights or activations, limiting deployment. Backdoor-based methods face additional pitfalls: ★ triggers built from high-perplexity, low-frequency tokens are easily caught by perplexity filters [34, 4]; ◼ fixed response patterns can be exposed by heuristic searches [14]; and ◆ spurious activations on benign inputs risk revealing the fingerprint and degrading usability [28, 34, 35, 46].
@@ 4.4 Model-Level Attack Robustness
Model Merging. We evaluate fingerprint identifiability when a fingerprinted model $\mathcal{M}_{\text{FP}}$ is merged with a donor $\mathcal{M}_{\text{D}}$ to form $\mathcal{M}_{\text{S}}$. Concretely on Mistral, we first apply unlearning to Mistral-7B-v0.3 to obtain the fingerprinted/unlearned model (i.e., $\mathcal{M}_{\text{FP}}$ = unlearned Mistral-7B-v0.3), and then merge it with Mistral-7B-Instruct-v0.3333https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 as the donor (i.e., $\mathcal{M}_{\text{D}}$ = Mistral-7B-Instruct-v0.3) using MergeKit [13]. We instantiate four strategies—Task, DARE-Task, TIE, and DARE-Tie—and for each we sweep the mixing ratio $\alpha\in\{0.1,0.2,\ldots,0.9\}$ (following MergeGuard [8]). For brevity, Table 4 reports the Task and TIE cases; the DARE variants exhibit the same trend. As donor weight increases, Chain&Hash degrades rapidly, IF-SFT remains partially resilient (rare-token triggers persist), while ForgetMark sustains high FSR across broad ratios across all strategies—suggesting that forgetting-based traces are intrinsically more robust than fixed trigger–response fingerprints.

### v2

@@ Abstract
footnotetext: *Equal contribution. Corresponding author.footnotetext: © 2026 IEEE. Published in ICASSP 2026 – 2026 IEEE International Conference on Acoustics, Speech and Signal Processing (ICASSP), scheduled for 3–8 May 2026 in Barcelona, Spain. Personal use of this material is permitted. However, permission to reprint/republish this material for advertising or promotional purposes or for creating new collective works for resale or redistribution to servers or lists, or to reuse any copyrighted component of this work in other works, must be obtained from the IEEE. Contact: Manager, Copyrights and Permissions / IEEE Service Center / 445 Hoes Lane / P.O. Box 1331 / Piscataway, NJ 08855-1331, USA. Telephone: +1 908 562 3966.
@@ 1 Introduction
Large language models (LLMs) and related paradigms such as vision–language models [21, 55, 53], autonomous agents  [39, 52, 54], and AI systems for security applications [17, 20, 18, 41, 35, 48] have become central to modern machine learning research and deployment, yet their creation depends on substantial proprietary data, compute, and engineering. Illicit model exfiltration and unauthorized fine-tuning have lowered the barrier to redistributing derivatives, creating urgent needs for reliable provenance auditing and copyright protection.
Model fingerprinting [40] has emerged as a practical remedy. Existing approaches fall into two families: intrinsic fingerprints, which rely on internal parameters or representations [5, 49, 47, 50], and invasive (backdoor-based) fingerprints, which implant a trigger–response mechanism detectable under black-box access [31, 45, 37, 4, 42, 46]. Intrinsic methods often require access to weights or activations, limiting deployment. Backdoor-based methods face additional pitfalls: ★ triggers built from high-perplexity, low-frequency tokens are easily caught by perplexity filters [37, 4]; ◼ fixed response patterns can be exposed by heuristic searches [14]; and ◆ spurious activations on benign inputs risk revealing the fingerprint and degrading usability [31, 37, 38, 51].
@@ 4.4 Model-Level Attack Robustness
Model Merging. We evaluate fingerprint identifiability when a fingerprinted model $\mathcal{M}_{\text{FP}}$ is merged with a donor $\mathcal{M}_{\text{D}}$ to form $\mathcal{M}_{\text{S}}$. Concretely on Mistral, we first apply unlearning to Mistral-7B-v0.3 to obtain the fingerprinted/unlearned model (i.e., $\mathcal{M}_{\text{FP}}$ = unlearned Mistral-7B-v0.3), and then merge it with Mistral-7B-Instruct-v0.3 [25] as the donor (i.e., $\mathcal{M}_{\text{D}}$ = Mistral-7B-Instruct-v0.3) using MergeKit [13]. We instantiate four strategies—Task, DARE-Task, TIE, and DARE-Tie—and for each we sweep the mixing ratio $\alpha\in\{0.1,0.2,\ldots,0.9\}$ (following MergeGuard [8]). For brevity, Table 4 reports the Task and TIE cases; the DARE variants exhibit the same trend. As donor weight increases, Chain&Hash degrades rapidly, IF-SFT remains partially resilient (rare-token triggers persist), while ForgetMark sustains high FSR across broad ratios across all strategies—suggesting that forgetting-based traces are intrinsically more robust than fixed trigger–response fingerprints.